# Substrate Scope Contrastive Learning: Repurposing Human Bias to Learn Atomic Representations

### v1

@@ Abstract
Learning molecular representation is a critical step in molecular machine learning that significantly influences modeling success, particularly in data-scarce situations. The concept of broadly pre-training neural networks has advanced fields such as computer vision, natural language processing, and protein engineering. However, similar approaches for small organic molecules have not achieved comparable success. In this work, we introduce a novel pre-training strategy, substrate scope contrastive learning, which learns atomic representations tailored to chemical reactivity. This method considers the grouping of substrates and their yields in published substrate scope tables as a measure of their similarity or dissimilarity in terms of chemical reactivity. We focus on 20,798 aryl halides in the CAS Content CollectionTM spanning thousands of publications to learn a representation of aryl halide reactivity. We validate our pre-training approach through both intuitive visualizations and comparisons to traditional reactivity descriptors and physical organic chemistry principles. The versatility of these embeddings is further evidenced in their application to yield prediction, regioselectivity prediction, and the diverse selection of new substrates. This work not only presents a chemistry-tailored neural network pre-training strategy to learn reactivity-aligned atomic representations, but also marks a first-of-its-kind approach to benefit from the human bias in substrate scope design.
@@ 1 Introduction
Encoding molecular structures into appropriate numerical representations that are computer-readable is key to accurate prediction of molecules’ reaction behavior, which is a pivotal challenge in chemical science and engineering. Traditionally, explaining reactivity based on chemical structure has relied on physical modeling and mechanistic analysis, along with the computation of key physical descriptors to build statistical models1, 2, 3, 4. This “feature engineering” approach relies on a prior understanding of relevant (computable) properties, which vary across different classes of molecules and reactions. In contrast, the advent of machine learning (ML)5 and, notably, deep learning (DL)6, has facilitated end-to-end learning as an alternative to manual crafting of molecular features. In principle, DL models can extract task-specific information from chemical structures, forming numerical features known as representations or embeddings. While already applied in modeling reaction yields7, 8, 9, most DL models remain remarkably ineffective or inaccurate in low data regimes.
The paradigm of pre-training deep neural networks on tasks with abundant data and adapting them to other tasks10, 11, 12, 13, 14 offers an attractive approach to enhance efficiency and generalizability. Ideally, such strategies enable few-shot learning, where models can learn new concepts from just a handful of examples15, 16. Successes in this approach are evident across computer vision 17, 18, 19, 20, natural language processing 21, 22, 23, and protein engineering24, 25. Inspired by these successes, numerous network pre-training strategies for molecular structures have been developed26, 27, 28, 29, 30. However, unlike in other domains, these methods have not yet led to substantial improvements in modeling 31.
We believe that there is a fundamental misalignment between existing pre-training tasks and the information crucial for downstream applications. Pre-training tasks for graph neural networks, for example, have involved predicting atom or functional group identities given the rest of the chemical structure,26, 32 which primarily teaches models about valence rules and the prevalence of certain functional groups. Similarly, pre-training tasks for models operating on SMILES strings33 teach models about similar principles along with the specifics of the SMILES language.34, 35, 36, 37 Recently introduced pre-training tasks involving the reconstruction of three-dimensional (3D) conformations, add steric considerations to the spectrum of information a model can learn 28, 30, 38, which represents progress but is still coarse for many of the chemical properties we are interested in.
Anticipating chemical reactivity 39—whether quantitatively (e.g., predicting yields) or qualitatively (e.g., predicting regioselectivity)—is a challenge where an ideal pre-trained molecular representation would incorporate information beyond valence rules, substructure popularity, and accessible conformations. To address this challenge, we introduce substrate scope contrastive learning (ContraScope), a new pre-training approach based on contrastive learning40, 41, 42 that leverages published reactions and benefits from human bias in substrate scope design (see Figure 1B and more statistics in Figure S2). The core idea of our method is that molecules exhibiting similar reactivity under identical circumstances should be considered similar in numerical representation as well. Given that one molecule can react in many disparate ways, we focus our representation learning goals on obtaining atom-level embeddings of specific reactive sites, rather than molecule-level embeddings. As a measure of synthetic similarity, we utilize substrate scope tables—collections of substrates, typically numbering 10-30, with their reaction yields reported under a consistent set of conditions. Published substrate scope tables often reflect a bias towards successful outcomes43. However, this bias suggests that, on average, two molecules within the same scope exhibit more similar reactivity than two molecules from different scopes. ContraScope is devised based on this principle.
@@
We demonstrate the efficacy of our contrastive learning approach using aryl halides, a key class of substrates employed in a wide range of chemical reactions. Specifically, we curated substrate scope tables from the literature published from 2010 to 2015, as indexed in the CAS Content CollectionTM, the largest human-curated collection of scientific data in the world, and demonstrate that our pre-training strategy successfully learns representations that capture key aspects of aryl halide reactivity. The learned embeddings show quantitative correlations with traditional reactivity descriptors and can be directly applied to tasks with limited data, such as yield prediction, regioselectivity prediction, and substrate selection. Our results support the hypothesis that substrate scope tables and publication bias, though previously considered detrimental to model training, actually offer valuable insights into reactivity patterns and can serve as a significant source of information. We hope to encourage the development of additional functionality-focused pre-training tasks—and pre-training tasks that effectively utilize human bias—to further the pursuit of universal molecular representations in chemistry.
@@ 2 Substrate Scope Contrastive Learning
Figure 1: (A) Substrate scope tables for training the network are curated from the CAS Content Collection, focusing on aryl halides with their associated yields. The substrate scope $i$, $j$ shown in the figure are real samples from the database44, 45. (B) Overview of substrate scope contrastive learning: Our pre-training strategy leverages the selective reporting bias in chemical reactions, partially revealed by the pronounced inclination towards higher yields in the histogram displaying the yield distribution for all surveyed reactions. Embeddings of substrates from the same scope table are pulled together, while embeddings of substrates from distinct scope tables are forced apart. A message-passing neural network operates on molecular graphs to derive atomic embeddings. The training aims to provide reactivity-aligned atomic embeddings learned from the substrate scopes, constructing a vector space representing reactivity trends. (C) The uniformly pre-trained embeddings can benefit various reactivity-related tasks.
The goal of our pre-training approach is to derive a multidimensional vector representation for the reactive atom in an aryl halide substrate that captures essential aspects of its reactivity. Ideally, this could be accomplished through supervised learning using a vast and varied dataset of molecules and their respective yields across numerous reaction conditions. However, the reality of data acquisition presents us with substrate scope tables that represent a narrow selection of molecules, influenced by the selective reporting tendencies and norms in scientific research and publication (see Figure S2 for statistics of the dataset). The core idea of substrate scope contrastive learning (ContraScope) is to use molecules not listed in a scope table as negative samples, leveraging the artificial bias in the substrate scope selection.
The model is trained on triplet of molecules: for each molecule as an anchor, we randomly select another molecule from the same substrate scope to serve as a positive sample, ensuring that their embeddings are closely aligned; concurrently, we sample a molecule that doesn’t belong to the substrate scope to act as a negative sample, ensuring their embeddings are distantly separated. Recognizing that molecules within the same substrate scope can exhibit varying degrees of reactivity, we define the ideal distance between the anchor and positive molecules as the difference in their yields. Formally, we train our model by minimizing the following loss function:
@@
Substrate scope tables and associated reaction yields are sourced from a subset of the CAS Content Collection describing publications and patents between 2010 and 2015. We restrict our analysis to aryl halides as they are an important category of molecules widely employed as building blocks in medicinal chemistry 46, 47. To ensure consistency within each substrate scope, we verified that all reactions involved transformations at a C–X bond and were performed under identical conditions, including reactants besides the aryl halides in the case of multi-component reactions, according to the database. We excluded scopes with fewer than five reactions, resulting in a total of 6,919 scopes with 64,192 reactions. For each training epoch, each aryl halide serves as the anchor in 16 triplets, which also include one randomly-sampled positive instance from the same scope and one randomly-sampled negative instance from a different scope. We adopt a graph isomorphism network (GIN)48, a widely used graph neural network, as the model architecture. The aryl halides are represented in graphs with atom and bond identities as input features49. Details on our training methodology are in the Methods Section.
@@ 3 Results
3.1 Learned aryl halide embeddings exhibit alignment with qualitative and quantitative measures of reactivity
@@
Figure 2: Qualitative analysis of aryl halide embeddings obtained through substrate scope contrastive learning. (A) A two-dimensional projection of learned embeddings using t-SNE50; each point denotes an aryl halide and is colored by halide type. Neighborhoods of similar aryl halides in the embedding space are annotated with structures, with the relevant C–X carbon highlighted. (B-C) Heatmap of the pair-wise distance between 12 para-substituted aryl bromides, showcasing that the learned embeddings’ signal-to-noise (SNR) distances (left) better reflects functional similarity than the conventional Tanimoto distances based on structural fingerprints (right).
@@
To evaluate the consistency of the learned aryl halide representations with human understanding of reactivity, we perform a combination of qualitative and quantitative analyses. We first visualize the learned embeddings to elucidate how the model has learned to organize the chemical space of aryl halides. To this end, we encoded a set of molecules comprising the 500 most frequently used aryl halides, augmented by a random sample of 2922 aryl halides from the substrate scope dataset. Upon encoding, we used t-distributed stochastic neighbor embedding (t-SNE) 50 projection to visualize the 64-dimensional embedding learned by the model in Figure 2A (see Figure S8 for results of other projection methods and a comparison with embeddings from an untrained network). A detailed examination of the position of specific structures in the embedding space (referred to as call-outs in Figure 2A) reveals that molecules with qualitatively similar reactivities—characterized by either electron-withdrawing groups (e.g., nitro, aldehyde, carbonyl) or electron-donating groups (e.g., hydroxyl, ether, amine, alkyl)—tend to cluster together. This pattern is consistent across various halide classes, underscoring the broad applicability of the embeddings we have developed.
In order to validate whether quantitative distances in the learned embedding space accurately reflect known functional similarity in chemical reactivity, we next encoded a set of 12 hand-selected para-substituted bromobenzenes, then calculated and visualized their pairwise distances (Figure 2B). The heatmap reveals that molecules with electron-withdrawing groups, such as nitrile, ester, and trifluoromethyl, demonstrate notably reduced distances compared to those with electron-donating groups like methoxy, hydroxyl, ether, and amine. This discernible pattern is not mirrored in the Tanimoto distances51, 52 (refer to Figure 2C), where distances are measured based on shared structural features in Morgan fingerprints. This comparison emphasizes that distances in our learned embedding space more effectively capture functional similarities in chemical reactivity than standard measures of structural similarity.
Figure 3: Relationship between the learned embeddings and conventional reactivity descriptors. (A) t-SNE visualizations of the learned embeddings, colored by traditional reactivity indicator values. (B) Regression performance ($r^{2}$) when learning to predict various physical organic chemistry descriptors from different input representations using linear models; negative values are excluded from the plot. Our embeddings, ContraScope, exhibit enhanced correlation with calculated reactivity descriptors compared to other common features. (C) Analysis of support vector machine (SVM)53 regression performance as a function of dataset size, highlighting the efficacy of our embeddings in scenarios with limited data.
@@
To further probe the correlation between our data-driven, substrate scope-informed embeddings and chemical reactivity, we embarked on a comparative analysis employing established reactivity descriptors2 calculated with first-principle quantum calculations54. We analyze a set of 762 aryl halides comprising of 500 of the most frequently appearing aryl halides from our dataset and 262 monosubstituted aryl halides with functional groups sourced from a Hammett constant table55. We calculated reactivity descriptors using Gaussian 1656 and auto-qchem57, with a particular focus on local atomic properties at the aromatic carbon bonded to the halogen, as our contrastive learning approach is designed to learn reactivity localized to the C–X motif, rather than molecule-level properties. Subsequent t-SNE projection mapping of this aryl halide set, as illustrated in Figure 3A, revealed that molecules positioned closely in the embedding space exhibited similar values of these atom-level reactivity descriptors. While the strongest differentiator is the identity of the halogen, more subtle patterns within each halide class can be seen.
To provide a more quantitative perspective of consistency, we trained simple (linear) machine learning models to predict various reactivity descriptors using the ContraScope embeddings as input features. We compare model performance using our embeddings, widely-used molecular descriptors58 or fingerprints59, and embeddings from two other popular pre-trained deep learning models. One is the same GIN architecture but pre-trained with a common attribute masking task (AttrMask)26, and another is a state-of-the-art chemical language model (ChemGPT)37. The regression performance of embeddings learned through substrate scope contrastive learning was comparable to, and often exceeded, these other representations (3B). As expected, these correlations are weaker for molecule-level features (Figure S16). Moreover, we also observe a stronger ability to predict reactivity descriptors starting from our embeddings in data-limited regimes (Figure 3C), highlighting the effectiveness of our pre-training strategy.
While these descriptors are all computable and do not truly need to be predicted from other molecular representations, these results illustrate a proof-of-concept for fine-tuning on reactivity-related tasks. The success of our purely data-driven approach, rooted in substrate scope selection, underscores the overlooked value of the human bias inherent in substrate scope tables, which unlocks a new information source for chemical reactivity modeling.
3.2 Pre-trained embeddings help enable various downstream applications
@@
Figure 4: Validation and application of learned embeddings in various downstream tasks. (A) Application of our embeddings to predict the yield of aryl bromides in cross-coupling reactions3. We show its predictive performance via leave-one-out validation compared with other common featurizations and a scatter plot mapping our predictions against experimental yields. (B) Application of our embeddings to predict the regioselectivity in arylation reactions of fluorobenzenes, with the model trained on penta- and trifluoronitrobenzenes to anticipate reactivity in tetrafluoronitrobenzenes, highlighting the reaction centers confirmed by experiments and annotated with prediction outcomes60. (C) Application of our learned embeddings for substrate scope selection, illustrated by a t-SNE plot representing purchasable aryl bromides categorized by clustering, with the chosen scope marked by cross symbols and detailed on the side.
The primary motivation for pursuing better molecular representations through pre-training is to enable downstream applications, particularly in low data regimes. We therefore select three example use cases with which to illustrate the promise of substrate scope contrastive learning.
@@
Yield prediction. As a first case study, we examine the prediction of reaction yields of various aryl bromides under identical reaction conditions. We select a Ni/photoredox catalyzed cross-coupling reaction, with reported yields for a diverse set of substrates published after 20153, and therefore not contained in our pre-training set. Kariofillis et al. report achieving a validation $r^{2}$ of $0.57$ with a univariate model based on a computed DFT descriptor, specifically electronegativity. Without being trained on any physical organic chemistry concepts or features (like electronegativity) and without requiring the cost of electronic structure calculations, our learned ContraScope embedding is able to achieve a comparable $r^{2}$ of 0.51 (Figure 4A). Other common fixed or pre-trained representations based on structure are far less successful.
@@
Regioselectivity prediction. As a second case study, we applied our embeddings to compare likelihood of multiple reactive sites within a molecule. We focused on a palladium-catalyzed Suzuki-Miyaura coupling reaction using polyfluoronitrobenzenes 60. The model was trained with yield data from penta- and trifluoronitrobenzenes, and used to predict the reactivity of different C-F sites in tetrafluoronitrobenzenes. The results, qualitatively confirmed by experimental data, show that a support vector machine (SVM) model using our aryl halide representations correctly learns to predict the expected reactivity at the C–F site ortho to the nitro group, but also distinguishes subtle differences in electronic effects, particularly where two adjacent C–F sites are nonequivalent (Figure 4B). This result highlights the proficiency of our embeddings in providing atomic-level representations to distinguish similar reactive sites and underscores the importance of learning atomic embeddings rather than molecular embeddings for reactivity prediction tasks.
Substrate scope selection. As our final case study, we illustrate how learned embeddings can assist in the selection of a “diverse” set of aryl bromides, which one may wish to do at the outset of an experimental screening campaign. This process, inspired by Kutchukian et al.’s61 and Kariofillis et al.’s3 methodology, involved clustering the chemical space of commercially-available aryl bromides using K-means62; while the authors used a DFT-derived feature vector, here we use our learned embeddings. A representative aryl bromide is chosen from each cluster to form a set designed to exhibit a diverse range of reactivities (Figure 4C). Unlike clustering based on structural fingerprints or expert-selected descriptors, our approach relies on the specific reactivity profiles of aryl halides as exemplified by our embeddings, in principle leading to a more curated set for this class of compounds better aligned with their unique reactivity characteristics. As there is no objective correct answer for diverse substrate selection, we leave the selection for qualitative interpretation by the reader.

### v2

# Revealing the Relationship Between Publication Bias and Chemical Reactivity with Contrastive Learning

@@ Abstract
A synthetic method’s substrate tolerance and generality are often showcased in a “substrate scope” table. However, substrate selection exhibits a frequently discussed publication bias: unsuccessful experiments or low-yielding results are rarely reported. In this work, we explore more deeply the relationship between such publication bias and chemical reactivity beyond the simple analysis of yield distributions using a novel neural network training strategy, substrate scope contrastive learning. By treating reported substrates as positive samples and non-reported substrates as negative samples, our contrastive learning strategy teaches a model to group molecules within a numerical embedding space, based on historical trends in published substrate scope tables. Training on 20,798 aryl halides in the CAS Content CollectionTM, spanning thousands of publications from 2010-2015, we demonstrate that the learned embeddings exhibit a correlation with physical organic reactivity descriptors through both intuitive visualizations and quantitative regression analyses. Additionally, these embeddings are applicable to various reaction modeling tasks like yield prediction and regioselectivity prediction, underscoring the potential to use historical reaction data as a pre-training task. This work not only presents a chemistry-specific machine learning training strategy to learn from literature data in a new way, but also represents a unique approach to uncover trends in chemical reactivity reflected by trends in substrate selection in publications.
@@ 1 Introduction
The report of a new reaction is concurrent with experiments showcasing the scope and limitations of the reaction. These insights are typically derived from evaluating a diverse array of substrates, typically under a constant or similar set of conditions, with results presented in substrate scope tables. However, the information presented in literature-reported substrate scope tables exhibits biases (see Figure 1B, with additional statistics in Figure S2) and is subject to various sociological factors 1, 2, 3, 4. This bias arises from two primary sources: selection bias, where chemists may choose substrates that are readily available or expected to give favorable results, and reporting bias, where high-yielding “successful” outcomes are more likely to be highlighted over low-yielding “negative” outcomes. These two factors interact synergistically, collectively termed publication bias2. While publication bias is pervasive across scientific disciplines5, its study within the context of chemistry, especially its relationship with chemical reactivity, remains in its early stages3.
Machine learning (ML) models are increasingly employed to model chemical reactivity quantitatively6, 7, 8, 9, opening up the possibility of studying reactivity beyond reported substrates. However, publication bias has led to a disproportionately low amount of negative data, making it difficult for ML models to reliably generalize to unseen compounds10, 3. In this paper, we approach the problem from a different perspective: we design a new training strategy leveraging the publication bias in substrate scope design as a premise, train a machine learning model that naturally reflects these historical trends, and examine its correlation with chemical reactivity. Specifically, we introduce substrate scope contrastive learning (ContraScope), a training strategy whose core premise is that substrates reported together in a substrate scope table (with similar yields) are more similar in reactivity than those not reported together on average. Using contrastive learning,11, 12, 13 we train a graph neural network model to encode reactive atoms into numerical embeddings, ensuring positive data are grouped closely while being distinguished from negative data. To account for variations within the same substrate scope table, we adjust the embedding distances among positive samples based on their reported yields under consistent conditions.
@@
We demonstrate our contrastive learning approach using aryl halides, a popular class of substrates employed in a wide range of chemical reactions. Specifically, we curated substrate scope tables from the literature published from 2010 to 2015, as indexed in the CAS Content CollectionTM, the largest human-curated collection of scientific data in the world, and demonstrate that the pattern of substrate usage in publications actually captures key aspects of aryl halide reactivity. The learned embeddings show both qualitative and quantitative correlations with traditional reactivity descriptors. Our results show that the publication bias in substrate scope tables, though previously considered detrimental to model training, actually offers valuable insights into reactivity patterns and can serve as a source of information. We hope to encourage further study of the relationship between publication bias and chemical reactivity, as well as the development of additional training strategies that effectively capture chemical reactivity—to further the pursuit of quantitative data-driven modeling in synthetic organic chemistry.
@@ 2 Substrate Scope Contrastive Learning
Figure 1: (A) Substrate scope tables for training the network are curated from the CAS Content Collection, focusing on aryl halides with their associated yields. The two substrate scopes $i$ and $j$ shown in the figure are real samples from the database14, 15. (B) Overview of substrate scope contrastive learning: The hypothesis behind our training strategy is that publication bias in chemical reactions reveals more subtle reactivity trends than the pronounced inclination towards higher yields as depicted in the histogram. A message-passing neural network operates on molecular graphs to derive atomic embeddings. Embeddings of substrates from the same scope table are pulled together, while embeddings of substrates from distinct scope tables are forced apart. The training aims to provide atomic embeddings that inherently reflect the grouping of reported substrate scopes table, which creates a vector space showcased to be aligned with reactivity trends. (C) The embeddings obtained after contrastive learning can then be analyzed to understand what has been learned and/or used as representations of molecules in various reactivity-related tasks.
The core idea of substrate scope contrastive learning (ContraScope) is to treat substrates reported within the same scope tables as positive samples while randomly sampling molecules that are not listed in that published substrate scope table as negative samples. By training the model this way, the resulting embeddings naturally reflect the selective design of substrate scope tables, and we can then investigate them to understand what trends have been learned. We emphasize that this approach does not assume every molecule excluded from a published scope table is incompatible with that reaction; rather, it uses the grouping of literature-reported substrate scope tables as a premise and source of inspiration. If there is some degree of publication bias–i.e., if the excluded substrates are, from a probabilistic perspective, less reactive or compatible than those included–then our learned embeddings should capture this underlying information about chemical reactivity.
Specifically, ContraScope trains the model on triplets of molecules: for each molecule as an anchor, we randomly select another molecule from the same substrate scope table to serve as a positive sample, encouraging their embeddings to be closely aligned; concurrently, we sample a molecule that doesn’t belong to the reported scope table to act as a negative sample, encouraging their embeddings to be distantly separated. Recognizing that molecules within the same reported substrate scope table can exhibit varying degrees of reactivity, we further encourage the distance between the anchor and positive molecules to be proportional to the difference in their yields. Formally, we train our model by minimizing the following loss function:
@@
Substrate scope tables and associated reaction yields are sourced from a subset of the CAS Content Collection describing publications and patents between 2010 and 2015. We restrict our analysis to aryl halides as they are an important category of molecules widely employed as building blocks in medicinal chemistry 16, 17. Within each substrate scope, we verified that all reactions involve transformations at a C–X bond and are performed under identical conditions, including reactants besides the aryl halides in the case of multi-component reactions, according to the database. We excluded scopes with fewer than five reactions, resulting in a total of 6,919 scopes with 64,192 reactions. For each training epoch, each aryl halide serves as the anchor in 16 triplets. We adopt a graph isomorphism network (GIN)18, a widely used graph neural network, as the model architecture. Taking a single substrate at a time as input, the model represents aryl halides as graphs with atom and bond identities as input features19. Details on our training methodology are in the Methods Section.
@@ 3 Results
3.1 Visualization and intuitive investigation of the learned aryl halide embeddings
@@
Figure 2: (A) A two-dimensional projection of learned embeddings using t-SNE20 for intuitive investigation of aryl halide embeddings obtained through substrate scope contrastive learning; each point denotes an aryl halide and is colored by halide type. Neighborhoods of similar aryl halides in the embedding space are annotated with structures. (B) t-SNE visualizations of the learned embeddings, colored by traditional reactivity descriptor values.
@@
We first visualize the learned embeddings to illustrate how the model has learned to organize the chemical space of aryl halides. To do this, we encode a set of molecules comprising the 500 most frequently used aryl halides, augmented by a random sample of 2922 aryl halides from the substrate scope dataset. Upon encoding, we use t-distributed stochastic neighbor embedding (t-SNE) 20 projection to visualize the 64-dimensional embedding learned by the model in Figure 2A (see Figure S16 for results of other projection methods and a comparison with embeddings from an untrained network).
A closer look at the position of specific structures in the embedding space (referred to as call-outs in Figure 2A) reveals that molecules with qualitatively similar reactivities—characterized by either electron-withdrawing groups (e.g., nitro, aldehyde, carbonyl) or electron-donating groups (e.g., hydroxyl, ether, amine, alkyl)—tend to cluster together. This pattern is consistent across various halide classes, underscoring the broad applicability of the embeddings we have developed. The pairwise distances of embeddings of 12 para-substituted bromobenzenes are visualized in Figure S15, revealing closer clustering for electron-withdrawing groups (e.g., nitrile, ester) compared to electron-donating groups (e.g., methoxy, amine). This trend, absent in Tanimoto distances from Morgan fingerprints (Figure S15), underscores the embedding space’s superior ability to capture functional reactivity similarities.
@@
To further investigate the connection between our data-driven, substrate scope-informed embeddings and chemical reactivity, we calculated and visualized established reactivity descriptors7 calculated with first-principle quantum calculations21 in the embedding space. We encode a set of 762 aryl halides comprising of 500 of the most frequently appearing aryl halides from our dataset and 262 monosubstituted aryl halides with functional groups sourced from a Hammett constant table22. We calculated reactivity descriptors using Gaussian 1623 and auto-qchem24, with a particular focus on local atomic properties at the aromatic carbon bonded to the halogen, as our contrastive learning approach is designed to learn reactivity localized to the C–X motif, rather than molecule-level properties. Subsequent t-SNE projection mapping of this aryl halide set, as illustrated in Figure 2B, revealed that molecules positioned closely in the embedding space exhibited similar values of these atom-level reactivity descriptors. While the strongest differentiator is the identity of the halogen, more subtle patterns within each halide class can be seen.
3.2 Regression analysis reveals the relationship between the learned embeddings and reactivity descriptors
Figure 3: Relationship between the learned embeddings and conventional reactivity descriptors. (A and B) Regression performance ($r^{2}$) when learning to predict local (B) and global (C) physical organic chemistry descriptors from different input representations using linear models; negative values are excluded from the plot. (C) Analysis of support vector machine (SVM)25 regression performance as a function of dataset size, highlighting the efficacy of our embeddings in scenarios with limited data.
To quantitatively assess how our learned embeddings capture chemical reactivity, we performed a regression analysis using simple (linear) machine learning models to predict various computed reactivity descriptors from the ContraScope embeddings. This approach allows us to evaluate the correlation between the learned pattern in our multidimensional embedding and descriptors we know to be relevant to chemical reactivity. By restricting our approach to simple linear models, we limit the overall modeling capacity; thus, strong regression performance would be driven primarily by the intrinsic relationship between the embedding features and the target descriptors. This would therefore gauge the extent to which publication trends alone reveal trends in chemical reactivity. We compare model performance using our embeddings, widely-used molecular featurizations26 or fingerprints27, as well as embeddings from three representative pre-trained deep learning models: (1) the same GIN architecture pre-trained using attribute masking (AttrMask)28, (2) a model pre-trained to capture three-dimensional molecular conformations (3DInfoMax)29, and (3) a state-of-the-art chemical language model (ChemGPT)30.
We calculated commonly used reactivity descriptors for the 762 aryl halides discussed in the previous section and performed regression analysis on them. Our learned embeddings outperformed or at least be comparable to other featurization and pre-trained embeddings in predicting most local reactivity descriptors related to the C–X motif at the carbon atom (Figure 3A), with the clearest advantages observed for partial charge-related descriptors and buried volume. Notably, as shown in Figure 3C, this performance gap becomes more significant when fewer data points are available, demonstrating that our pre-training procedure provides informed embeddings as a robust initialization that markedly improves downstream modeling performance, indicating the close relationship between learned representations and both Mulliken charges and buried volumes. As the training set size increases, we observe the apparent advantage in representation to deteriorate and expect the advantage to disappear entirely at even larger training set sizes. In contrast, correlations were weaker for global (molecule-level) reactivity descriptors (Figure 3B), which is expected due to the design of our pre-training strategy focusing on reactivity localized to the C–X motif. Overall, these findings confirm that the multi-dimensional patterns learned from the historical trend in substrate scope grouping strongly correlate with local reactivity characteristics while exhibiting a lower correlation with global reactivity descriptors.
3.3 Potential downstream application of the learned embeddings
@@
Figure 4: Validation and application of learned embeddings in various downstream tasks. (A) Application of ContraScope embeddings to predict the yield of aryl bromides in cross-coupling reactions8. We show its predictive performance via leave-one-out validation compared with other common featurizations and a scatter plot mapping our predictions against experimental yields. (B) Application of ContraScope embeddings to predict the regioselectivity in arylation reactions of fluorobenzenes, with the model trained on penta- and trifluoronitrobenzenes to anticipate reactivity in tetrafluoronitrobenzenes, highlighting the reaction centers confirmed by experiments and annotated with prediction outcomes31.
In this section, we showcase potential downstream applications of our learned embeddings for reaction performance prediction tasks. These evaluations serve to demonstrate how information relevant to the downstream tasks is implicitly revealed by publication bias. If it were not, we would expect ContraScope embeddings to be unable to perform these tasks. We do not aim to or claim to achieve state-of-the-art performance on supervised learning tasks of reaction performance.
@@
Yield prediction. As a first case study, we examine the prediction of reaction yields of various aryl bromides under identical reaction conditions. We select a Ni/photoredox catalyzed cross-coupling reaction, with reported yields for a diverse set of substrates published after 20158, and therefore not contained in our pre-training set. Kariofillis et al. report achieving a validation $r^{2}$ of $0.57$ with a univariate model based on a computed DFT descriptor, specifically electronegativity. Without being trained on any physical organic chemistry concepts or features (like electronegativity) and without requiring the cost of electronic structure calculations, our learned ContraScope embedding is able to achieve a comparable $r^{2}$ of 0.51 (Figure 4A). Other common fixed or pre-trained representations based on structure are far less successful, further confirming that trends in substrate scope publication contain information relevant to reactivity.
@@
Regioselectivity prediction. As a second case study, we applied our embeddings to compare the likelihood of multiple reactive sites within a molecule. We focused on a palladium-catalyzed Suzuki-Miyaura coupling reaction using polyfluoronitrobenzenes 31. The model was trained with yield data from four penta- and trifluoronitrobenzenes, and used to predict the reactivity of different C-F sites in three tetrafluoronitrobenzenes. The results, quantitatively confirmed by experimental data, show that a support vector machine (SVM) model using our aryl halide representations correctly learns to predict the expected reactivity at the C–F site ortho to the nitro group, but also distinguishes subtle differences in electronic effects, particularly where two adjacent C–F sites are nonequivalent (Figure 4B). This result highlights that ContraScope’s atomic-level representations can distinguish similar reactive sites and underscores the importance of learning atomic embeddings rather than molecular embeddings for reactivity prediction tasks.